OpenStudy (anonymous):

what is the product of the 2 solutions of the equation x^2+3x-21=0

OpenStudy (anonymous):

imagine alpha and beta is two solutions of the equation ax^2 + bx + x = 0 so: alpha = {-b + sqr(b^2 - 4ac)} \ 2 and beta = {-b - sqr(b^2 - 4ac)}

OpenStudy (anonymous):

SRY! alpha = {-b + sqr(b^2 - 4ac)} / 2a and beta = {-b - sqr(b^2 - 4ac)} / 2a

OpenStudy (anonymous):

so product would be c/a and c=-21 and a=1 so it would be -21

OpenStudy (ahsome):

Step 1: Find the two x-intercepts of the equation Step 2: Multiply the x-intercepts together X-intercepts can be found using the quadratic formula: \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Sub in your values: \[\frac{-1*3\pm\sqrt{3^2-4*1*-21}}{2*1}\] Find the two values. In this case, it was: \(-6.32, 3.32\) Multiply the numbers together: \(-6.32*3.32=-20.9824\)

OpenStudy (anonymous):

@Ahsome , no need to calculate b^2 - 4ac ;) look at my solution.

OpenStudy (anonymous):

the product of the solutions of the equation ax^2 + bx + c = 0 is c/a

OpenStudy (skullpatrol):

Any questions @babygirl65 ?

OpenStudy (anonymous):

Do you know how to solve this equation because im not sure the one above is right @skullpatrol

OpenStudy (zzr0ck3r):

@PFEH.1999 `\(\alpha=\frac{-b+\sqrt{b^2-4ac}}{2a}\)` will give \(\alpha=\frac{-b+\sqrt{b^2-4ac}}{2a}\)

OpenStudy (skullpatrol):

Yes, I do. What does the question ask for @babygirl65 ?

OpenStudy (anonymous):

what is the product of the 2 solutions of the equation x^2+3x-21=0 @skullpatrol

OpenStudy (skullpatrol):

Yes the answer to that question can be seen by knowing that in the form x^2 + bx + c the number c is the product of the two solutions

OpenStudy (anonymous):

thank you

OpenStudy (skullpatrol):

Thanks for asking :-)

OpenStudy (perl):

kudos to all the responses

OpenStudy (zzr0ck3r):

what @PFEH.1999 said is great for the form \(ax^2+bx+c=0\) if we let \(\alpha\) and \(\beta\) be our two solutions we know the answers will be \(\alpha=\frac{-b+\sqrt{b^2-4ac}}{2a}\) and \(\beta=\frac{-b-\sqrt{b^2-4ac}}{2a}\) then \(\alpha*\beta=( \frac{-b-\sqrt{b^2-4ac}}{2a})*(\frac{-b+\sqrt{b^2-4ac}}{2a})= \frac{(-b)^2-b^2+4ac}{4a^2}=\frac{c}{a}\)

OpenStudy (skullpatrol):

but @zzr0ck3r we don't need that form here :/

OpenStudy (zzr0ck3r):

ha sorry was not trying to byte your style...

OpenStudy (zzr0ck3r):

what do you mean you don't need that from here? this gives the general solution to any quadratics.

OpenStudy (zzr0ck3r):

as long as the determinant is non negative...

OpenStudy (skullpatrol):

All I'm saying is why not solve this question as a simpler special case of the general formula, for the sake of clarity @zzr0ck3r

OpenStudy (zzr0ck3r):

you just asked a mathematician that question:P

OpenStudy (skullpatrol):

:O

OpenStudy (zzr0ck3r):

it was a very clever solution, I only typed it out because I thought the original poster was not great it latex, but I was wrong and waster my time.

OpenStudy (ahsome):

Wow. That answer is GENIOUS. WOAH